A collection of dimes and nickles is worth $8.40. There are 106 coins in all. How many of each?
Answer by checkley77(12844)

N+D=106
N=106-D
.05N+/10D=8.40
.05(106-D)+.10D=8.40
5.3-.05D+.10D=8.40
.05D=8.40-5.30
.05D=3.10
D=3.10/.05
D=62 DIMES.
106-62=44 NICKLES.
PROOF:
.05*44+.10*62=8.40
2.20+6.20=8.40
8.40=8.40